OpenStudy (anonymous):

How do I factor x^4+81 I know that x=3i but I need the factoring. Any help is welcomed. :)

OpenStudy (mathstudent55):

I assume you are familiar with: \(a^2 = b^2 = (a + b)(a - b)\)

OpenStudy (mathstudent55):

\(x^4 + 81\) \(= x^4 - (-81)\) \(= (x^2)^2 - (9i)^2\) \(= (x ^2 + 9i)(x^2 - 9i)\)

OpenStudy (anonymous):

Oh thank you so much. I have been dealing with figuring too much out. This helps greatly!

OpenStudy (mathstudent55):

It's not done, though. Now it gets more complicated.

OpenStudy (anonymous):

That is fine I just needed the start. :) I am having to solve with the imaginary and real parts. If I run into trouble I may hit you up. ;)